### 04 Mar 2026 v1

# MOOSE-Star: Unlocking Tractable Training for Scientific Discovery by Breaking the Complexity Barrier

Abstract
While large language models (LLMs) show promise in scientific discovery, existing research focuses on inference or feedback-driven training, leaving the direct modeling of the generative reasoning process, $P(\text{hypothesis}|\text{background})$ ($P(h|b)$), unexplored. We demonstrate that directly training $P(h|b)$ is mathematically intractable due to the combinatorial complexity ($O(N^{k})$) inherent in retrieving and composing inspirations from a vast knowledge base. To break this barrier, we introduce MOOSE-Star, a unified framework enabling tractable training and scalable inference. In the best case, MOOSE-Star reduces complexity from exponential to logarithmic ($O(\log N)$) by (1) training on decomposed subtasks derived from the probabilistic equation of discovery, (2) employing motivation-guided hierarchical search to enable logarithmic retrieval and prune irrelevant subspaces, and (3) utilizing bounded composition for robustness against retrieval noise. To facilitate this, we release TOMATO-Star, a dataset of 108,717 decomposed papers (38,400 GPU hours) for training. Furthermore, we show that while brute-force sampling hits a “complexity wall,” MOOSE-Star exhibits continuous test-time scaling.
Machine Learning, ICML
1 Introduction
Recently, increasing research attention has been brought to the field of LLMs for scientific discovery. However, most research efforts to date focus on developing inference methodologies to leverage LLMs for discovery, with only a few works investigating how to train an LLM specifically for this task (Luo et al., 2025). Critically, nearly all of these existing training approaches rely on external feedback to refine generated hypotheses, rather than explicitly modeling the core conditional probability $P(\text{hypothesis}|\text{background})$ (denoted as $P(h|b)$). In essence, these methods learn how to update a hypothesis given feedback, but overlook the fundamental reasoning process required to generate a high-quality hypothesis directly from the research background. For example, Weng et al. (2024); Li et al. (2024) leverage open-access reviews to train critique models; Behzadifar et al. (2025) utilize alignment with observed data as reward; and Goel et al. (2025) employ rubric-based LLM feedback.
Consequently, how to train $P(h|b)$ remains unknown. In this paper, we provide a theoretical analysis demonstrating why directly training $P(h|b)$ is intractable due to its inherent combinatorial complexity. Furthermore, we propose a recipe to overcome this barrier, enabling the tractable and, crucially, scalable training of $P(h|b)$.
To understand this intractability, we build upon the formalization established in prior studies (Yang et al., 2025b; Liu et al., 2025), viewing the generation of a hypothesis $h$ as a composition of the research background $b$ and a sequence of latent inspirations retrieved from the global knowledge base. For instance, the discovery of Backpropagation can be viewed as composing the background of “multilayer logistic regression” with the inspiration of the “chain rule” from calculus. This compositional view is supported by cognitive science studies (Koestler, 1964; Benedek et al., 2012) and recent inference-based discovery methods (Yang et al., 2024; Wang et al., 2024b; Garikaparthi et al., 2025; Kumar et al., 2025). However, directly training $P(h|b)$ implies implicitly learning to retrieve the correct sequence of $k$ inspirations from a literature space of size $N$ (e.g., $N\approx 10^{7}$, encoded within the LLM’s pre-trained parameters). This results in a combinatorial search space of $O(N^{k})$, rendering end-to-end training mathematically intractable.
To break this complexity barrier, we introduce MOOSE-Star (MS), a framework that transforms the intractable objective into a solvable form through a series of theoretical decompositions. Specifically, we leverage the mathematical formulation proposed by Yang et al. (2025b), which decomposes the combinatorial complexity of $P(h|b)$ into a set of subtasks with linear complexity: retrieving inspirations (sequentially) and composing each inspiration into a hypothesis. Crucially, while Yang et al. (2025b) utilized this equation solely for inference, we operationalize it here for training. This decouples the exponential $O(N^{k})$ dependency into $k$ independent sequential steps, reducing the sample complexity to a linear $O(k\times N)$.
However, even with this decomposition, explicitly scanning the full literature ($N$) remains computationally prohibitive. To address this, we introduce three key innovations: (1) Hierarchical Search: We organize the global literature into a semantic search tree, replacing the expensive linear scan with top-down navigation. This reduces the retrieval complexity from linear $O(N)$ to logarithmic $O(\log N)$ in the best-case scenario. (2) Bounded Composition: To handle retrieval imperfections, we introduce a semantic tolerance radius. By training the composition module on noisy inspirations within a “bounded window” centered on the exact inspiration $i^{*}$, we ensure the model is robust to inexact retrievals. (3) Motivation Planning: We demonstrate that explicitly modeling a “Motivation” variable—appended to the research background—acts as a dynamic generative root for the search tree. By effectively pruning irrelevant branches during hierarchical search, this reduces the number of search steps required to locate the relevant semantic subspace ($N_{m}<N$).
To support this data-hungry training paradigm, we developed a comprehensive pipeline that processes scientific literature across domains such as biology, chemistry, and cognitive science. In total, we extracted research backgrounds, ground-truth hypotheses, and inspirations—each explicitly linked to a historical citation—from 108,717 papers, a massive computational effort consuming approximately 38,400 A800 GPU hours. We release this processed dataset as the TOMATO-Star dataset.
Finally, we demonstrate in § 7 that MOOSE-Star achieves superior test-time scaling compared to brute-force sampling. While brute-force sampling hits a complexity wall on hypotheses requiring the composition of multiple inspirations, our inspiration-based hierarchical search exhibits a continuous improvement in success rate with increased inference budget, effectively turning an intractable discovery problem into a manageable search process.
Our contributions can be summarized as below:
  • We establish the first theoretical analysis on why training $P(h|b)$ is intractable (combinatorial complexity).
  • We provide the first training recipe on how to enable the tractable and scalable training of $P(h|b)$.
  • We provide an inference recipe that supports more scalable test-time inference for discovery.
  • We release a dataset of 108,717 processed papers (cost 38,400 GPU hours), along with the full training & inference codebase and trained models.
2 Related Work
While a significant body of research focuses on novel inference algorithms for LLM-based discovery (Romera-Paredes et al., 2024; Novikov et al., 2025; Shojaee et al., 2025; Lu et al., 2024), few studies investigate methods for training LLMs for discovery. Most existing training efforts focus on constructing external feedback mechanisms for generated hypotheses rather than modeling the core conditional probability $P(h|b)$ itself. For instance, Weng et al. (2024); Li et al. (2024) leverage peer review data to train critique models, which serve as reward signals for the discovery agent. Similarly, in the domain of equation discovery, Behzadifar et al. (2025) derive rewards based on the goodness of fit between the proposed hypothesis (equation) and observed data. Goel et al. (2025) extract evaluation criteria from scientific literature to construct rubric-based LLM self-evaluators, while Pu et al. (2025) train models to retrieve ancillary high-level guidance from a constrained set to assist discovery. O’Neill et al. (2025) attempt to model $P(h|b)$ directly; however, obtaining the reasoning trace towards $h$ from only $b$ via distillation is prohibitive (as discussed in § 6), and generating simulated reasoning traces directly from $b$ and the ground truth $h$ as input is very difficult to work (Wang et al., 2025).
This work focuses exclusively on training LLMs for scientific discovery, distinct from related fields such as general scientific reasoning (Bai et al., 2025) or modeling of specific science modality (Jumper et al., 2021).
3 Preliminary: The Decomposition Theory
Our framework builds upon the probabilistic decomposition theory introduced in MOOSE-Chem (Yang et al., 2025b). Fundamentally, this theory posits that a scientific hypothesis $h$ is not generated ex nihilo (out of nothing) but is synthesized by composing a research background $b$ with a sequence of $k$ latent inspirations, $\mathbf{i}=(i_{1},\ldots,i_{k})$, retrieved from a global knowledge base $\mathcal{I}$ ($N=|\mathcal{I}|$). Formally, the generation of a hypothesis can be modeled as:
$h=f(b,i_{1},\ldots,i_{k})$ (1)
This compositional view is well-supported by cognitive science literature (Koestler, 1964) and has proven empirically effective across scientific disciplines (Liu et al., 2025). For instance, the discovery of Backpropagation can be viewed as composing the background of “multilayer logistic regression” with the inspiration of the “chain rule” from calculus; similarly, a new synthesis method in chemistry might be derived by applying a specific “catalyst” (inspiration) to a known reaction mechanism.
Based on this formulation, we can approximate the intractable marginal likelihood $P(h\mid b)$ by decomposing the generation process into $k$ sequential steps. Assuming a Markov property where the intermediate hypothesis $h_{j-1}$ sufficiently encapsulates the history of prior inspirations, the probability can be factorized using the chain rule as:
$P(h\mid b)\approx\prod_{j=1}^{k}\underbrace{P(i_{j}\mid b,h_{j-1},\mathcal{I})}_{\text{Inspiration Retrieval}}\cdot\underbrace{P(h_{j}\mid b,h_{j-1},i_{j})}_{\text{Hypothesis Composition}}$ (2)
where $h_{j}$ denotes the intermediate hypothesis state after the $j$-th step (with $h_{0}=b$ and $h_{k}=h$). The first term represents the retrieval of the next relevant inspiration, and the second term represents the composition of the hypothesis using the retrieved inspiration.
4 Dataset Construction
To enable the training of our framework, we construct TOMATO-Star, a large-scale dataset of scientific papers processed into structured components. The construction pipeline consists of four stages: collection, decomposition, structured representation, and quality assurance.
Data Collection and Split. We collected 108,717 open-access papers from the NCBI database (Schoch et al., 2020), spanning biology, chemistry, and cognitive science. The corpus covers the period from January 2020 to October 2025. To rigorously evaluate generalization, we enforce a strict temporal split: papers published from January 2020 to September 2025 constitute the training set, while papers from October 2025 serve as the held-out test set. We also perform explicit checking to make sure the two sets do not have overlap with each other. October 2025 is latest enough, so that the evaluation results should not be contaminated.
Preprocessing and Decomposition. Raw PDF documents are first converted to Markdown format using MinerU (Wang et al., 2024a). We then employ a pipeline of locally deployed reasoning models, specifically DeepSeek-R1 and R1-distilled-Qwen-32b (Guo et al., 2025), to decompose each paper into a tuple of $(b,h,\mathbf{i})$: Research Background, Hypothesis, and Inspirations. This workflow adapts the design of ResearchBench (Liu et al., 2025) but introduces stricter quality controls and a novel hypothesis representation scheme.
Structured Representation. The decomposed components are defined as follows:
  • Research Background ($b$): Comprises the specific research question and a background survey summarizing prior methods addressing that question.
  • Inspirations ($\mathbf{i}$): Ground-truth inspirations are identified directly from the source paper’s citations. We augment these citations by retrieving their full titles and abstracts via Semantic Scholar (Kinney et al., 2023).
  • Hypothesis ($h$) as Incremental Deltas: Unlike standard summaries, we structure the hypothesis as a sequence of “Delta Hypotheses” ($\Delta h$), enforcing a one-to-one mapping where each inspiration $i_{j}$ results in exactly one $\Delta h_{j}$. The final hypothesis is the concatenation of these deltas. Crucially, each $\Delta h$ is structured into three levels: (1) Motivation (why this direction was chosen), (2) Mechanism (why it can work), and (3) Methodology (how it is implemented).
Quality Assurance. To ensure data integrity, every processed sample must pass four automated quality checks prior to inclusion: (1) Information Necessity: Each proposed $i$ must provide essential, complementary knowledge required to derive $h$ from $b$. (2) Information Sufficiency: The integration of the background and inspirations must logically entail the hypothesis ($b+\mathbf{i}\approx h$). (3) Information Disjointness: The background $b$ must remain strictly independent, leaking no information present in either the inspirations or the hypothesis. (4) Non-Redundancy: The extracted inspirations within $\mathbf{i}$ must be mutually distinct and non-repetitive.
5 Methodology
In this section, we first analyze why directly training $P(h|b)$ is computationally intractable. We then derive the Moose-Star framework, showing how each component progressively transforms the intractable objective into a solvable problem with manageable complexity.
5.1 Analysis: The Intractability Barrier ($O(N^{k})$)
Consider the standard baseline approach: training a model to directly maximize the marginal likelihood $P(h\mid b)$ end-to-end. As established in Eq. 1, generating a valid hypothesis inherently requires identifying and synthesizing a sequence of $k$ inspirations $\{i_{1},\ldots,i_{k}\}$. Consequently, direct modeling $P(h\mid b)$ implies an implicit search over the Cartesian product of the knowledge base, $\mathcal{I}^{k}$. Even disregarding the computational cost of composition, the search space complexity alone scales exponentially:
$\mathcal{C}_{\text{E2E}}>O(N^{k})\quad(\text{e.g., for }N=10^{7},k=3:\approx 10^{21})$ (3)
Here, $N=|\mathcal{I}|\approx 10^{7}$ represents the scale of global scientific literature, where each paper constitutes an independent unit of knowledge. Crucially, this complexity barrier persists even in the absence of an external retriever. When relying solely on internal weights, the LLM is forced to implicitly retrieve from the knowledge encoded during pre-training. Given this intractable $O(N^{k})$ search space, standard end-to-end training becomes mathematically ill-posed, resulting in severe convergence difficulties.
5.2 Method $\mathrm{I}$: Decomposed Sequential Training ($O(N)$)
To dismantle the exponential barrier of $O(N^{k})$, our first methodology shifts from directly training the monolithic $P(h\mid b)$ to translating the theoretical decomposition in Eq. 2 into a concrete training objective. We train the model on two equivalent sequential subtasks: Inspiration Retrieval and Hypothesis Composition.
For each step $j$, the complexity breakdown is as follows:
  • Inspiration Retrieval (IR): The task is to identify the correct ground-truth inspiration from the full global knowledge base. This requires scanning the search space of size $N$, resulting in a complexity of $O(N)$.
  • Hypothesis Composition (HC): The task is to generate $h_{j}$ given the retrieved inspiration $i_{j}$. Conditioned on the provision of the exact ground-truth inspiration, the generation in the ideal case requires only a single inference pass. Thus, relative to the search space size $N$, this step has a constant complexity of $O(1)$.
Since this process is repeated sequentially for $k$ steps, this methodology fundamentally transforms the search space topology from a Cartesian product ($N^{k}$) to a linear sum ($k\times N$). Consequently, the complexity is:
$\mathcal{C}_{\mathrm{I}}\approx k\times\Big[\underbrace{O(N)}_{\text{Inspiration Retrieval}}+\underbrace{O(1)}_{\text{Hypothesis Composition}}\Big]$ (4)
We formulate both IR and HC as generative reasoning tasks, requiring the model to produce a Chain-of-Thought (CoT) (Wei et al., 2022) rationale before outputting the final answer. To optimize this reasoning capability, we adopt teacher-based Rejection Sampling Fine-Tuning (RFT) (Touvron et al., 2023; Zelikman et al., 2022). Specifically, we utilize R1-Distilled-Qwen-32B as the teacher model to generate training samples, which are then filtered to fine-tune the student model, R1-Distilled-Qwen-7B.
For the IR task, the model is trained to identify the target inspiration via a generative selection process. The input consists of the query context and a candidate pool of 15 papers (1 positive and 14 negatives), each represented by its title and abstract. The negative set comprises easy negatives (randomly sampled papers published no later than the source paper) and hard negatives drawn from three specific categories: papers with keyword overlap with the source background, papers with keyword overlap with the ground-truth inspiration, and papers semantically similar to the ground-truth inspiration. We acknowledge that these negatives might occasionally indicate valid alternative inspirations. However, the ground-truth citation represents the author’s explicit acknowledgment of causal influence, providing the most reliable supervision signal for prioritizing optimal inspirations over merely related ones. Furthermore, this single-selection objective minimizes the penalty for potentially valid unlabelled candidates compared to independently labeling each candidate.
For the HC task, the objective is to generate the incremental hypothesis update, $\Delta h_{j}$. The model is conditioned on the research background $b$, the retrieved ground-truth inspiration $i_{j}$, and $h_{j-1}$, represented by the concatenation of $\{\Delta h_{0},\ldots,\Delta h_{j-1}\}$ (if $j\geq 1$). To enable effective rejection sampling, we developed a rubric-based evaluator to compare the quality of sampled hypotheses against the ground-truth $\Delta h_{j}$, ensuring that only high-quality samples are retained for training. The evaluation rubric is provided in § C.
5.3 Method $\mathrm{II}$: Bounded Composition
While the linear decomposition $O(kN)$ breaks the exponential barrier, linearly scanning $N\approx 10^{7}$ papers for every retrieval step remains computationally prohibitive for test-time inference. To address this, we must further optimize the linear term $O(N)$.
We achieve this by relaxing the retrieval objective. In § 5.2, the formulation assumes a strict “Exact Match” constraint: the IR module must identify the unique ground-truth inspiration $i^{*}$ (11 out of $N$), and the HC module relies solely on this exact input. We propose Bounded Composition, which introduces a Semantic Tolerance Space of size $M$. Instead of requiring $i^{*}$, we posit that the HC module can be trained to function robustly within a semantic neighborhood $\mathcal{I}_{i^{*}}$ centered on $i^{*}$ containing $M$ candidates (where $i^{*}\in\mathcal{I}_{i^{*}}\subset\mathcal{I}$). If the retriever provides any inspiration $i\in\mathcal{I}_{i^{*}}$, the HC module is expected to reason through this proxy $i$ to recover the underlying $i^{*}$ and correctly compose the hypothesis.
This relaxation alters the complexity landscape:
  • IR Relaxation ($O(N)\to O(N/M)$): The retrieval target expands from a single point to a subspace of size $M$. In a randomized search or sampling process, the expected number of steps to hit a target of size $M$ in a space of $N$ reduces linearly to $N/M$.
  • HC Overhead ($O(1)\to O(M)$): The composition cost increases. The HC model must now perform additional reasoning to resolve the ambiguity within the tolerance space. We represent this abstractly as an internal complexity of $O(M)$, reflecting the cost of reasoning within the local space of $M$ possibilities.
Consequently, the overall complexity updates to:
$\mathcal{C}_{\{\mathrm{I},\mathrm{II}\}}\approx k\times\Big[\underbrace{O\left(\frac{N}{M}\right)}_{\text{Relaxed Retrieval}}+\underbrace{O(M)}_{\text{Bounded Composition}}\Big]$ (5)
This trade-off is strategic: while HC complexity increases linearly with $M$, the dominant IR complexity decreases by a factor of $M$. Given that $N\gg M$, shifting the computational burden from the global search space to local reasoning is a significant net reduction in total complexity.
Figure 1: Bounded Composition. The concentric circles around $i^{*}$ represent similarity thresholds that define the bounded space $M$.
To operationalize Bounded Composition, we first curate a pool of semantic proxies. For each ground-truth inspiration $i^{*}$ in the dataset, we query Semantic Scholar to retrieve up to 50 semantically similar papers. We embed these candidates using SPECTER2 (Singh et al., 2022) and compute their cosine similarity to $i^{*}$. Based on these scores, we stratify valid candidates into three difficulty tiers: Easy ($[0.94,0.97)$), Medium ($[0.92,0.94)$), and Hard ($[0.90,0.92)$). From each available tier, we select the single candidate with the highest similarity score to serve as the tier representative. We then employ the teacher model to generate hypotheses conditioned on these proxy inspirations rather than the exact ground truth. The generated outputs are rigorously evaluated against the quality control rubric described in § C. Crucially, for instances where proxies from multiple tiers yield passing hypotheses, we prioritize the sample from the most difficult tier (i.e., lowest similarity) for the final training set. This strategy explicitly maximizes the learned semantic tolerance space, forcing the model to reason robustly even with less precise retrieval contexts.
5.4 Method $\mathrm{III}$: Hierarchical Search ($O(\log N)$)
Despite the reduction to $O(N/M)$, a linear scan over the global knowledge base $\mathcal{I}$ remains inefficient. To enable logarithmic-time access (at least in the best-case scenario), we propose replacing the flat scan with a Hierarchical Best-First Search. This design draws inspiration from MOOSE-Chem2 (Yang et al., 2025a), which formulates fine-grained hypothesis discovery as a hierarchical search over varying levels of detail abstraction. Here, we adapt this concept by constructing the hierarchy directly over the inspiration space $\mathcal{I}$. By organizing $\mathcal{I}$ into a semantic search tree, we can rapidly navigate to the most promising subspaces, pruning irrelevant branches early in the process.
Tree Construction (Offline).
We construct the search tree via a bottom-up recursive clustering approach. First, all papers in $\mathcal{I}$ are embedded using SPECTER2. We then perform hierarchical K-means clustering (McQueen, 1967) recursively from the leaves up to the root. In each iteration, we cluster the current level’s embeddings (initially the papers, subsequently the cluster centroids), where the resulting centroids serve as the nodes for the next higher level of the hierarchy. This process repeats until a single root centroid remains. Crucially, to align with our IR model’s capacity, we enforce a maximum branching factor of $c=15$. We employ a post-clustering balancing operation: if a cluster exceeds size $c$, its furthest member (least similar to the centroid) is reassigned to its nearest centroid in the same hierarchy that has available capacity (size <c<c). This ensures the tree is balanced with a consistent branching factor.
Best-First Search (Online).
During inference, we traverse the tree using a Best-First Search (BFS) strategy. We maintain a priority queue of candidate nodes, ranked by a path confidence score. In each iteration, we pop the highest-scoring node for expansion. The IR model evaluates the sub-branches of this node, generating a probability distribution over its children (derived from the output token log-probabilities). Crucially, these local probabilities are aggregated with the parent node’s stored cumulative probability to compute updated scores for each child, which are then pushed back into the priority queue. To ensure fair comparison between nodes at different depths—avoiding the bias against deeper paths inherent in raw cumulative products—we define the ranking score as the geometric mean of the probabilities along the path from the root:
$\text{Score}(path_{j})=\sqrt[j+1]{\prod_{t=0}^{j}p_{t}}$ (6)
where $p_{t}$ is the probability assigned by the IR model at depth $t$. By dynamically updating the search frontier based on this length-normalized metric, the algorithm effectively prioritizes the most promising semantic neighborhoods without traversing the entire database.
In the best-case scenario (where the IR model makes ideal routing decisions), the search complexity corresponds to the depth of the tree, which is logarithmic with respect to the search space. This yields more tractable complexity:
$\mathcal{C}_{\{\mathrm{I},\mathrm{II},\mathrm{III}\}}\approx k\times\Big[\underbrace{O\left(\log\frac{N}{M}\right)}_{\text{Hierarchical Search}}+\underbrace{O(M)}_{\text{Bounded Composition}}\Big]$ (7)
5.5 Method $\mathrm{IV}$: Motivation Planning
While hierarchical search optimizes the traversal of the knowledge base, it is essentially an unguided navigation: the search process often lacks explicit direction, essentially “stumbling” into candidates without prior intent. To address this, we introduce a Motivation variable ($m$).
Semantic Guidance.
Functionally, $m$ serves as a Directional Guide generated prior to retrieval. By appending $m$ to the query context $b$, we explicitly condition the hierarchical search on specific intended properties or requirements. This biases the search focus towards relevant semantic manifolds, filtering out branches inconsistent with the high-level intent. Consequently, the effective search space is reduced from the global knowledge base $N$ to a motivation-aligned subspace $N_{m}$ (where $N_{m}<N$).
O(1) Generation.
Crucially, the motivation $m$ is derived solely from the research background $b$ via internal abstraction (e.g., a single-step translation of background information into high-level requirements). We intentionally design this planning step to be lightweight—approximately $O(1)$ (a single inference)—to ensure it neither becomes a computational bottleneck nor generates overly specific directives that could misguide the subsequent computationally expensive IR and HC resources. Assuming the generated motivation is sound (a reasonable assumption for this lightweight translation step), the overall complexity is further reduced to:
$\mathcal{C}_{\text{MS}}\approx k\times\Big[\underbrace{O(1)}_{\text{Motivation}}+\underbrace{O\left(\log\frac{N_{m}}{M}\right)}_{\text{Focused Retrieval}}+\underbrace{O(M)}_{\text{Bounded Composition}}\Big]$ (8)
Hierarchical MDP Formulation.
This addition formally extends the decomposition in Eq. 2 into a tri-stage Hierarchical Markov Decision Process (HMDP). At each constructive iteration $j$, the model first formulates a high-level intent $m_{j}$ before executing retrieval and composition:
$\begin{split}P(h\mid b)\approx\prod_{j=1}^{k}\Big[&\underbrace{P(m_{j}\mid b,h_{j-1})}_{\text{Motivation Planning}}\cdot\underbrace{P(i_{j}\mid b,h_{j-1},m_{j},\mathcal{I})}_{\text{Inspiration Retrieval}}\\ &\cdot\underbrace{P(h_{j}\mid b,h_{j-1},m_{j},i_{j})}_{\text{Hypothesis Composition}}\Big]\end{split}$ (9)
The formal derivation of Eq. 9 is provided in § B. Compared to the base formulation, the additional term $P(m_{j}\mid b,h_{j-1})$ explicitly governs the motivation planning step, ensuring that the subsequent retrieval ($i_{j}$) and composition ($h_{j}$) remain aligned with a coherent research trajectory.
6 Experiment
In this section, we empirically evaluate the four methodologies proposed in § 5. All experiments utilize the TOMATO-Star test set, comprising papers published in October 2025, to ensure a rigorous evaluation free from contamination.
6.1 Method $\mathrm{I}$: Decomposed Sequential Training
Table 1: Inspiration retrieval results.
Model Accuracy
Random Selection 6.70%
R1-Distilled-Qwen-7B 28.42%
MS-IR-7B 54.37%
Table 2: Hypothesis composition performance given the ground-truth inspiration $i^{*}$. “Total” aggregates the scores for Motivation (Mot), Mechanism (Mec), and Methodology (Met). “w/ $n\times$ bounded” indicates the model was trained with $n$ additional copies of bounded composition data.
Model Total Mot Mec Met Length
R1-Distilled-Qwen-7B 4.34 1.96 1.40 0.97 231.02
MS-HC-7B 5.08 2.21 1.58 1.29 204.12
   w/ $1\times$ bounded 5.16 2.24 1.60 1.31 203.84
   w/ $2\times$ bounded 5.15 2.25 1.58 1.32 205.17
Inspiration Retrieval (IR).
We fine-tuned R1-Distilled-Qwen-7B on a dataset of 152k retrieval examples, resulting in MS-IR-7B (where MS stands for MOOSE-Star). As shown in Table 1, this specialized training yields a substantial performance gain, boosting accuracy from 28.42% to 54.37%.
Hypothesis Composition (HC).
We fine-tuned R1-Distilled-Qwen-7B on 97k samples generated via teacher-based rejection sampling, resulting in MS-HC-7B (MS stands for MOOSE-Star). Performance is evaluated using the M3 score (detailed in § C), an LLM-based rubric summing scores for Motivation, Mechanism, and Methodology. Each dimension ranges from [0, 4] (representing 0%–100% information recall), yielding a total score of 12.
Table 2 presents the results on ground-truth inputs. The fine-tuned MS-HC-7B significantly outperforms the baseline. Furthermore, results indicate that incorporating “bounded” training data (hypotheses generated from noisy inspirations) further enhances performance, suggesting that exposure to semantic variations during training improves reasoning robustness even when evaluating on perfect inspirations.
Table 3: Hypothesis composition performance under varying levels of inspiration noise. The tiers Easy (E), Medium (M), and Hard (H) correspond to the semantic similarity of the retrieved proxy inspiration to the ground truth. “w/ $n\times$ bounded” indicates the model was trained with $n$ additional copies of bounded composition data. Bold values denote the best performance in each column.
Easy (E) Medium (M) Hard (H)
Model Total Mot Mec Met Total Mot Mec Met Total Mot Mec Met
R1-Distilled-Qwen-7B 2.58 1.43 0.62 0.54 2.16 1.24 0.46 0.45 1.82 1.13 0.33 0.36
MS-HC-7B 2.82 1.53 0.62 0.67 2.37 1.36 0.47 0.55 2.20 1.30 0.40 0.49
   w/ $1\times$ bounded 2.96 1.55 0.70 0.71 2.52 1.41 0.53 0.58 2.19 1.25 0.44 0.49
   w/ $2\times$ bounded 3.03 1.62 0.69 0.72 2.49 1.37 0.53 0.59 2.25 1.29 0.44 0.52
6.2 Method $\mathrm{II}$: Bounded Composition
Table 3 evaluates performance under noisy conditions, using proxy inspirations stratified by semantic similarity to $i^{*}$ (Easy, Medium, Hard). The results demonstrate that incorporating bounded training data significantly boosts M3 scores across all tiers compared to the base model.
6.3 Method $\mathrm{III}$: Hierarchical Search
Table 4: Efficiency and accuracy of hierarchical search variants compared to a tournament baseline. IR inference calls: Average number of IR model queries required to identify the ground-truth inspiration. Proposed Rank: The average position of the ground-truth inspiration in the retrieved list (lower is better).
Search Method IR inference calls Proposed Rank
Tournament Search 218.00 987.76
Hierarchical Search 67.78 813.40
   w/ motivation (simple) 63.80 767.64
   w/ motivation (detailed) 63.05 742.50
We evaluate search efficiency on a corpus of 3,035 ground-truth inspirations derived from the 1,658 test papers (October 2025). A hierarchical tree is constructed over this corpus as described in § 5.4. Table 4 compares our Hierarchical Best-First Search against a Tournament Search baseline. The baseline employs an exhaustive bottom-up strategy, processing all candidates through the cluster structure to identify winners, which incurs a high, fixed computational cost. We report two metrics: IR Inference Calls (the average number of IR queries required to successfully locate $i^{*}$) and Proposed Rank (the average position of $i^{*}$ in the retrieval order, where lower is better).
The results demonstrate that our hierarchical approach is significantly more efficient, reducing inference calls by approximately $3\times$ (67.78 vs. 218.00) while achieving a superior (lower) average rank. This confirms that the top-down, probability-guided search effectively prunes irrelevant branches without compromising retrieval accuracy.
6.4 Method $\mathrm{IV}$: Motivation Planning
We analyze the impact of motivation quality on search efficiency by utilizing ground-truth derived directives. We compare two variants: (1) Simple: A direct translation of requirements from $b$ with minimal reasoning and specification. (2) Detailed: Derived from the delta hypothesis, which we observe typically captures strategic intent without leaking specific mechanism or methodology details.
As shown in Table 4, incorporating motivation consistently improves search efficiency over the baseline. The detailed variant achieves the best performance, confirming that more detailed and precise planning can reduce more computational effort required to locate $i^{*}$.
7 Scaling Analysis
7.1 Scaling Brute-force Sampling $P(h|b)$
Table 5: Comparison of sampling pass rates between Brute-force (End-to-End) and Hypothesis Composition (Decomposed). While Brute-force performance collapses to 0.00% as complexity ($k$) increases, the decomposed HC module maintains a robust 47.33% pass rate, unlocking the data engine for RFT.
Method Steps ($k$) Total Samples Passed Samples Pass Rate
Brute-force $P(h|b)$ 1 31,314 653 2.09%
2 67,000 90 0.13%
3 12,000 0 0.00%
Total 110,314 743 0.67%
HC $P(\Delta h|i,b)$ 1 (per step) 195,301 92,427 47.33%
Direct distillation from the end-to-end distribution $P(h|b)$ is intractable due to the extreme scarcity of valid reasoning traces. Table 5 illustrates this bottleneck: finding a qualified hypothesis is rare even for simple cases ($k=1$), and becomes statistically impossible for multi-inspiration trajectories ($k\geq 2$) due to combinatorial explosion.
In contrast, the decomposed HC step demonstrates a high success rate (47.33%). This performance gain is driven by two structural advantages. Primarily, by isolating a single inspiration step, the HC task circumvents the combinatorial complexity of the full problem: it is only required to solve for one incremental innovation $\Delta h_{j}$ rather than the joint distribution of all $k$ innovations. Secondarily, the decomposition outsources the search burden to the IR module; thus, conditioning on the provided inspiration $i^{*}$ explicitly guides the generation. Together, these factors empirically validate the complexity analysis in § 5.
Evaluation Standard and The “Training Deadlock”.
To understand why this performance gap determines training feasibility, we must examine the criteria used for harvesting RFT training data (detailed in § C). Both strategies use the same rubric, evaluating a generated hypothesis component against the ground truth across Motivation, Mechanism, and Methodology. To serve as a positive sample for RFT, we impose a filtering threshold: the generation must achieve a Total Score $\geq 8$ (out of 12).
The structural difference lies in how this threshold applies to the two paradigms:
  • Brute-force: Since the baseline generates the hypothesis $h$ holistically, a valid sample for a paper with $k\geq 2$ must successfully meet the filtering threshold for every constituent innovation $\Delta h_{j}^{*}$ ($j=1\dots k$) simultaneously. This strict conjunctive constraint drives the success rate to near-zero, creating a “training deadlock”: the model cannot generate enough valid samples to initiate RFT, rendering the training process impossible.
  • HC: In contrast, the HC module is trained to synthesize a single increment $\Delta h_{j}$. Crucially, leveraging the decomposed data construction, we can condition the model on the ground-truth inspiration $i_{j}^{*}$. Under this setup, the model only needs to meet the threshold for the corresponding single $\Delta h_{j}^{*}$. Consequently, this efficient sampling capability breaks the training deadlock, making it feasible to train HC using RFT.
7.2 Scaling Training of MOOSE-Star
We analyze the scaling behavior of the IR and HC models with training data ranging from $10^{0}$ to $10^{5}$ samples (Figure 2(a) and 2(b)). The IR model demonstrates a standard log-linear improvement in accuracy. In contrast, the HC model exhibits a distinct threshold behavior: significant log-linear performance gains emerge only after the dataset size exceeds $10^{3}$ samples. One potential reason for this disparity is that IR is fundamentally a classification task, whereas HC is a more complex generative task requiring higher data density to capture the nuances of scientific synthesis.
Scalability of Decomposition.
These results are pivotal as they validate the training paradigm established in Equation 2. They demonstrate that the decomposed subtasks—Inspiration Retrieval (IR) and Hypothesis Composition (HC)—accomplish more than simply converting the intractable problem of scientific discovery into tractable, learnable sub-problems; they also establish a scalable training paradigm that yields log-linear performance improvements with increasing data scale.
(a) Inspiration retrieval.
(b) Hypothesis composition.
Figure 2: Scaling laws of MOOSE-STAR: (a) IR and (b) HC.
Generalization on OOD Retrieval.
(a) Cumulative success vs inference steps.
(b) Win-rate by problem complexity.
Figure 3: Test-time scaling behavior of MOOSE-STAR versus brute-force.
The continuous scaling of the IR task is particularly noteworthy and counter-intuitive. Prior literature typically characterizes scientific inspiration retrieval as an Out-of-Distribution (OOD) task (Yang et al., 2025b), given that true discovery requires associating concepts that have never been linked in the prior knowledge base. Despite this inherent novelty barrier, the consistent log-linear improvement suggests that the model is not merely memorizing seen connections, but is potentially acquiring a generalizable “logic of discovery”—an abstract intuition for identifying meaningful scientific associations across previously unidentified knowledge connections.
7.3 Test-time Scaling of MOOSE-Star
Figure 3(a) compares the test-time scaling efficiency of MOOSE-Star against the brute-force baseline. The evaluation is conducted on a subset of TOMATO-Star comprising the first 109 test cases. This sample size corresponds to the first 200 sequential inspiration steps in the dataset, as each research paper in our corpus typically involves a composition of $k\in[1,3]$ inspirations.
The x-axis represents the computational cost measured in cumulative inference calls. For MOOSE-Star, this total sums (1) all IR inference calls required to navigate the hierarchical search tree, and (2) the number of HC inference calls equivalent to the proposed rank of the groundtruth inspiration. Specifically, for an inspiration $i^{*}$ proposed at rank $n$, we account for $n$ separate HC composition attempts: every retrieved inspiration candidate $i$ usually needs a HC model inference to compose hypothesis. For the baseline, the cost simply equals the total number of independent samples generated.
The y-axis tracks the success rate, defined as the proportion of cases where a generated hypothesis matches or exceeds the quality of the MOOSE-Star final output, judged by how well it captures the innovative essence of $h^{*}$. We utilize the MOOSE-Star output as the reference standard for two main reasons: (1) Direct side-by-side comparison is the most robust method for determining relative superiority between two approaches; (2) Since MOOSE-Star is conditioned on the complete set of ground-truth inspirations ($i^{*}$), it inherently serves as a high-quality proxy for the ground truth, setting a strong baseline for validating whether unguided methods can recover the key scientific concepts. We implement the pairwise comparison with Gemini-3-Flash (Google DeepMind, 2025). To mitigate the influence of position bias and inference variance, each pair is evaluated four times (across two independent trials, each including both original and swapped positions).
The results reveal a stark contrast in scaling behavior. MOOSE-Star achieves steady progress, continuously solving more complex cases as the search budget increases, eventually reaching 100% coverage of the test set at $\sim$6,000 inference calls. Conversely, the brute-force baseline saturates at $\sim$41.3%.
Figure 3(b) provides a breakdown of this saturation, revealing the steep “complexity cliff” where unstructured sampling fails. For MOOSE-Star, each HC inference is effectively “seeded” by a specific retrieved inspiration $i$, allowing for a diverse and directed exploration of the hypothesis space. In contrast, the baseline relies on unguided sampling. The impact of this difference is profound: even when fully penalized for search costs (by accounting for “proposed rank” in HC inference calls), MOOSE-Star essentially relies on a single guided hypothesis sample (conditioned on the retrieved $i^{*}$) to compete against a massive “barrage” of 9,500 independent brute-force samples.
On simple, single-step tasks ($k=1$), the baseline leverages this sheer volume to achieve a $\sim$53% success rate. However, this volume advantage is rapidly negated by the combinatorial explosion. As the reasoning chain extends, the baseline’s performance collapses—plummeting to $\sim$36% for $k=2$ and further evaporating to a negligible $\sim$8% for $k=3$. This dramatic decay empirically confirms that without the structural guidance of decomposition, even nearly ten thousand unguided attempts are statistically insufficient to break the combinatorial complexity barrier.
8 Conclusion
We address the intractability of training LLMs for scientific discovery by demonstrating that end-to-end modeling of $P(h|b)$ is ill-posed due to combinatorial complexity. To overcome this, we introduce MOOSE-Star, which operationalizes the probabilistic decomposition proposed in MOOSE-Chem to enable tractable training. Our results confirm that the synergy of Hierarchical Search, Bounded Composition, and Motivation Planning reduces search complexity from exponential to logarithmic in the best-case scenario. Crucially, while brute-force methods hit a “complexity wall” on hypotheses requiring multiple inspirations, MOOSE-Star exhibits continuous test-time scaling, transforming discovery into a manageable search process. We release TOMATO-Star to facilitate future research.
Appendix A MOOSE-Star Concept Figure
Figure 4: MOOSE-Star Concept Figure
Appendix B Theoretical Derivation: Rigorous Decomposition of $P(h\mid b)$ via Hierarchical MDP
B.1 Extended Fundamental Assumption
Building upon the foundational assumption in (Yang et al., 2025b) that scientific hypotheses originate from background knowledge and inspirations, we introduce the concept of Motivation ($m$). We posit that the discovery process is not merely a selection of entities, but a hierarchical decision process: before identifying a specific inspiration $i$ (e.g., a specific paper or entity), a researcher first formulates a high-level strategic intent or motivation $m$ (e.g., a methodology category or research direction).
We translate this extended assumption into the following functional form:
$h=f(b,(m_{1},i_{1}),\ldots,(m_{k},i_{k}))$ (10)
Here, $m_{j}$ serves as the semantic anchor or abstract strategy for the $j$-th step, and $i_{j}$ is the concrete inspiration instantiation conditioned on $m_{j}$. The tuple $(m_{j},i_{j})$ represents a composite unit of inspiration for each step.
Consistent with the previous theoretical framework (Yang et al., 2025b), we maintain two critical assumptions to ensure mathematical tractability by collapsing the otherwise intractable marginalization space:
  1. 1.
    Uniqueness Assumption: Each valid hypothesis $h$ corresponds to a unique minimal set of motivation-inspiration pairs, $\{(m_{1},i_{1}),\ldots,(m_{k},i_{k})\}$. Motivation: Instead of assuming a hypothesis is an arbitrary, intractable amalgamation of global knowledge, this posits that a discovery is driven by a specific, identifiable set of core insights. This allows us to concentrate the probability mass entirely on this single optimal set of ingredients.
  2. 2.
    Fixed-Order Assumption: The integration of this set follows a canonical constructive sequence (e.g., chronologically or by logical dependency). Motivation: In empirical sciences, the order in which inspirations are integrated can often be interchangeable. However, adopting a fixed canonical order serves to simplify the mathematical exposition by eliminating the need to marginalize over all $k!$ possible permutations of the selected pairs, effectively reducing the generation process to a single, deterministic reasoning path. (We provide the generalized formulation relaxing this assumption at the end of this section).
B.2 Hierarchical Markov Decision Process Formulation
We formalize the hypothesis generation process as a Hierarchical Markov Decision Process (Hierarchical MDP). Unlike a standard MDP where the action is a single step, we define the decision at step $j$ as a hierarchical process involving a high-level planner and a low-level executor:
  • State: $s_{j-1}=(b,h_{j-1})$, representing the background context and the intermediate hypothesis generated so far (with $h_{0}=\emptyset$).
  • High-level Action (Motivation): $m_{j}\in\mathcal{M}$, selected by a high-level policy $\pi_{high}$. This action abstracts the search space.
  • Low-level Action (Inspiration): $i_{j}\in\mathcal{I}$, selected by a low-level policy $\pi_{low}$. This action is conditioned on $m_{j}$.
  • Transition (Composition): The environment updates the state to $s_{j}=(b,h_{j})$ by integrating the selected inspiration $i_{j}$ into the previous intermediate hypothesis $h_{j-1}$ under the guidance of the motivation $m_{j}$.
B.3 Derivation of the Decomposition
Our goal is to decompose the intractable marginal likelihood $P(h\mid b)$. Let $I$ denote the deterministic global knowledge base. Since it serves as a fixed, omnipresent context for the discovery process, conditioning on the background $b$ is practically equivalent to conditioning on the joint context $(b,I)$.
Under the uniqueness and fixed-order assumptions, the marginalization over all possible latent variables collapses to the joint probability of the unique constructive path defined by the sequence $\{(m_{1},i_{1}),\ldots,(m_{k},i_{k})\}$:
$P(h\mid b)$ $=P(m_{1},i_{1},\ldots,m_{k},i_{k},h_{1},\ldots,h_{k}\mid b)$ (11)
$=\prod_{j=1}^{k}P(m_{j},i_{j},h_{j}\mid b,m_{1},i_{1},h_{1},\ldots,m_{j-1},i_{j-1},h_{j-1})$ (12)
$\approx\prod_{j=1}^{k}P(m_{j},i_{j},h_{j}\mid b,h_{j-1},I)$ (13)
Equation 11 establishes the joint distribution of the unique path. Equation 12 applies the standard chain rule over the time steps $j=1\dots k$, explicitly capturing the full historical trajectory up to step j1j-1. Equation 13 applies the Markov property, assuming the decision at step $j$ depends primarily on the research background $b$, the global knowledge base $I$ (introduced here as the explicit search space), and the immediate previous hypothesis state $h_{j-1}$.
Intra-step Decomposition.
Now, we perform the decomposition for the joint term at a single step $j$, $P(m_{j},i_{j},h_{j}\mid b,h_{j-1},I)$. We first apply the exact chain rule and then apply conditional independence assumptions inherent to the scientific discovery process:
$P(m_{j},i_{j},h_{j}\mid b,h_{j-1},I)$
$=P(m_{j}\mid b,h_{j-1},I)\cdot P(i_{j}\mid b,h_{j-1},m_{j},I)\cdot P(h_{j}\mid b,h_{j-1},m_{j},i_{j},I)$ (14)
$\approx\underbrace{P(m_{j}\mid b,h_{j-1})}_{\text{{Step 1: Planning}}}\cdot\underbrace{P(i_{j}\mid b,h_{j-1},m_{j},I)}_{\text{{Step 2: Retrieval}}}\cdot\underbrace{P(h_{j}\mid b,h_{j-1},m_{j},i_{j})}_{\text{{Step 3: Composition}}}$ (15)
The transition from Equation 14 to 15 is justified by two independence properties:
  1. 1.
    Internal Planning Independence: The strategic motivation $m_{j}$ is an internal cognitive state generated by the model’s reasoning over the context $(b,h_{j-1})$. It is conditionally independent of the external database $I$ given the context ($P(m\mid\dots,I)\approx P(m\mid\dots)$).
  2. 2.
    Local Composition Independence: Once a specific inspiration $i_{j}$ is retrieved, the formation of the next hypothesis state $h_{j}$ relies on the semantic content of $i_{j}$ and the current context. The rest of the global database $I$ becomes redundant ($P(h\mid\dots,i_{j},I)\approx P(h\mid\dots,i_{j})$).
Substituting Equation 15 back into Equation 13, we obtain the final tractable objective:
$P(h\mid b)\approx\prod_{j=1}^{k}\Big[\underbrace{P(m_{j}\mid b,h_{j-1})}_{\text{Motivation Planning}}\cdot\underbrace{P(i_{j}\mid b,h_{j-1},m_{j},I)}_{\text{Inspiration Retrieval}}\cdot\underbrace{P(h_{j}\mid b,h_{j-1},m_{j},i_{j})}_{\text{Bounded Composition}}\Big]$ (16)
Generalization without the Fixed-Order Assumption.
While the fixed-order assumption significantly simplifies the theoretical exposition, in practice, the components of a scientific hypothesis can often be assembled in various valid sequences. Relaxing this assumption requires marginalizing over all valid permutations $\Pi_{k}$ of the $k$ steps. The generalized decomposition for MOOSE-STAR is thus:
$P(h\mid b)$ $=\sum_{\pi\in\Pi_{k}}P(m_{\pi(1)},i_{\pi(1)},\ldots,m_{\pi(k)},i_{\pi(k)},h_{1}^{(\pi)},\ldots,h_{k}^{(\pi)}\mid b)$
$\approx\sum_{\pi\in\Pi_{k}}\prod_{j=1}^{k}\Big[P(m_{\pi(j)}\mid b,h_{j-1}^{(\pi)})\cdot P(i_{\pi(j)}\mid b,h_{j-1}^{(\pi)},m_{\pi(j)},I)\cdot P(h_{j}^{(\pi)}\mid b,h_{j-1}^{(\pi)},m_{\pi(j)},i_{\pi(j)})\Big]$ (17)
where $h_{0}^{(\pi)}=\emptyset$, $h_{k}^{(\pi)}=h$, and $h_{j}^{(\pi)}$ denotes the intermediate hypothesis state at step $j$ under the specific permutation $\pi$.
Appendix C Rubrics for Comparing Generated Hypothesis with Groundtruth Hypothesis
Scoring Rubric (0–4 for each dimension)
IMPORTANT INSTRUCTIONS:
  1. 1.
    Score based on RECALL – what percentage of GT content is correctly covered by Generated.
  2. 2.
    Both MISSING and WRONG content count as “not covered”.
  3. 3.
    The examples below are ONLY for illustration – do NOT match against them.
    Score the actual GT vs Generated content, not similarity to examples.
Motivation (WHY) – Does it identify the same research gap?
How to score: Count what percentage of GT’s key elements appear correctly in Generated.
[EXAMPLE FOR ILLUSTRATION ONLY – do not match against this]
Example GT:
Current deep learning methods for brain tumor segmentation in MRI scans suffer from low accuracy at tumor boundaries, particularly in low-contrast glioma regions, due to insufficient modeling of boundary uncertainty
(5 key elements: domain=brain tumor/MRI, task=segmentation, problem=low boundary accuracy, context=low-contrast glioma, cause=insufficient uncertainty modeling)
  • 4 ($\sim$100%):
    Existing DL approaches for brain tumor MRI segmentation have poor boundary delineation in glioma cases because they fail to model boundary uncertainty
    $\checkmark$ All 5: domain + task + problem + context + cause
  • 3 ($\sim$75%):
    Missing: “Brain tumor segmentation methods have low accuracy at boundaries in gliomas”
    $\checkmark$ 4: domain + task + problem + context  —  $\times$ Missing: cause
    Wrong: “Brain tumor segmentation in MRI has low boundary accuracy in gliomas due to limited training data”
    $\checkmark$ 4: domain + task + problem + context  —  $\times$ Wrong: cause (limited data vs uncertainty modeling)
  • 2 ($\sim$50%):
    Missing: “Brain tumor segmentation in MRI has accuracy issues”
    $\checkmark$ 2.5: domain + task + vague problem  —  $\times$ Missing: context + cause
    Wrong: “Brain tumor detection in MRI suffers from false positives in gliomas”
    $\checkmark$ 2: domain + context  —  $\times$ Wrong: task (detection) + problem (false positives)
  • 1 ($\sim$25%):
    Missing: “Medical image segmentation needs improvement”
    $\checkmark$ 1: broad domain only  —  $\times$ Missing: specific target + problem + context + cause
    Wrong: “Brain tumor classification methods are too slow”
    $\checkmark$ 1: target organ  —  $\times$ Wrong: task (classification) + problem (speed)
  • 0 ($\sim$0%):
    “Protein structure prediction lacks accuracy”
    $\checkmark$ 0  —  $\times$ Completely unrelated domain
    “Natural language models struggle with long-range dependencies”
    $\checkmark$ 0  —  $\times$ Completely unrelated domain
Mechanism (HOW IT WORKS) – Does it propose the same core mechanism?
GT:
Apply transformer-based attention with boundary-aware loss functions to learn multi-scale feature representations, enabling precise tumor boundary localization through uncertainty-guided refinement
(5 key elements: architecture=transformer attention, loss=boundary-aware, features=multi-scale, task=boundary localization, technique=uncertainty refinement)
  • 4 ($\sim$100%):
    Use transformer attention with boundary loss for multi-scale features to localize tumor boundaries via uncertainty-guided refinement
    $\checkmark$ All 5: architecture + loss + features + task + technique
  • 3 ($\sim$75%):
    Missing: “Transformer attention with boundary-aware loss for multi-scale feature learning to localize tumor boundaries”
    $\checkmark$ 4: architecture + loss + features + task  —  $\times$ Missing: refinement technique
    Wrong: “Transformer attention with boundary loss for multi-scale features to localize boundaries via post-processing CRF”
    $\checkmark$ 4: architecture + loss + features + task  —  $\times$ Wrong: technique (CRF vs uncertainty)
  • 2 ($\sim$50%):
    Missing: “Use attention mechanism with boundary loss for tumor boundary detection”
    $\checkmark$ 2.5: partial architecture + loss + task  —  $\times$ Missing: multi-scale + refinement
    Wrong: “Use transformer attention with standard loss for multi-scale feature learning”
    $\checkmark$ 2.5: architecture + features  —  $\times$ Wrong: loss (standard vs boundary) + missing: task + technique
  • 1 ($\sim$25%):
    Missing: “Apply deep learning for tumor analysis”
    $\checkmark$ 0.5: broad method category  —  $\times$ Missing: specific architecture + loss + features + technique
    Wrong: “Use transformer attention for image classification”
    $\checkmark$ 1: architecture  —  $\times$ Wrong: task (classification) + missing: loss + features + technique
  • 0 ($\sim$0%):
    “Apply LSTM for time series forecasting”
    $\checkmark$ 0  —  $\times$ Completely unrelated mechanism and task
    “Use rule-based heuristics for text classification”
    $\checkmark$ 0  —  $\times$ Completely unrelated mechanism and domain
Methodology (HOW IT’S INTEGRATED) – Does it describe similar implementation?
GT:
Train on BraTS 2021 dataset (1251 MRI cases), implement 3D U-Net with transformer encoder, use combined Dice-boundary loss, apply 5-fold cross-validation, evaluate with Dice score and 95% Hausdorff distance
(6 key details: dataset=BraTS 2021/1251, architecture=3D U-Net + transformer, loss=Dice-boundary, validation=5-fold CV, metrics=Dice + HD95)
  • 4 ($\sim$100%):
    Train on BraTS 2021 (1251 scans), 3D U-Net with transformer encoder, Dice-boundary loss, 5-fold CV, report Dice and HD95
    $\checkmark$ All 6: dataset + architecture + loss + validation + metrics
  • 3 ($\sim$75%):
    Missing: “BraTS 2021 dataset, 3D U-Net + transformer, Dice-boundary loss, 5-fold CV, Dice score”
    $\checkmark$ 5: dataset + architecture + loss + validation + partial metrics  —  $\times$ Missing: HD95
    Wrong: “BraTS 2021 (1251 scans), 3D U-Net + transformer, Dice-boundary loss, 3-fold CV, Dice and HD95”
    $\checkmark$ 5: dataset + architecture + loss + metrics  —  $\times$ Wrong: validation (3-fold vs 5-fold)
  • 2 ($\sim$50%):
    Missing: “Train on BraTS dataset with 3D U-Net, evaluate Dice score”
    $\checkmark$ 3: partial dataset + architecture + partial metrics  —  $\times$ Missing: size + loss + validation + HD95
    Wrong: “Train on private dataset, use 2D U-Net, Dice-boundary loss, 5-fold CV, Dice”
    $\checkmark$ 3: loss + validation + partial metrics  —  $\times$ Wrong: dataset + architecture (2D vs 3D)
  • 1 ($\sim$25%):
    Missing: “Train a segmentation model on brain MRI data”
    $\checkmark$ 1.5: vague dataset + vague approach  —  $\times$ Missing: specific details
    Wrong: “Train on TCGA genomic data, use ResNet, cross-entropy, accuracy”
    $\checkmark$ 0.5: general training  —  $\times$ Wrong: dataset + architecture + loss + metrics
  • 0 ($\sim$0%):
    “Fine-tune GPT on dialogue dataset with RLHF”
    $\checkmark$ 0  —  $\times$ Completely different domain and methodology
    “Survey 200 patients using questionnaires, analyze with chi-square test”
    $\checkmark$ 0  —  $\times$ Completely different methodology type

### 06 Apr 2026 v2

@@
 By effectively pruning irrelevant branches during hierarchical search, this reduces the number of search steps required to locate the relevant semantic subspace ($N_{m}<N$).
-To support this data-hungry training paradigm, we developed a comprehensive pipeline that processes scientific literature across domains such as biology, chemistry, and cognitive science.
+To support this data-hungry training paradigm, we developed a comprehensive pipeline that processes scientific literature across domains such as biology, chemistry, medicine, medical imaging, psychology, and cognitive science.
 In total, we extracted research backgrounds, ground-truth hypotheses, and inspirations—each explicitly linked to a historical citation—from 108,717 papers, a massive computational effort consuming approximately 38,400 A800 GPU hours.
@@
 While brute-force sampling hits a complexity wall on hypotheses requiring the composition of multiple inspirations, our inspiration-based hierarchical search exhibits a continuous improvement in success rate with increased inference budget, effectively turning an intractable discovery problem into a manageable search process.
-Our contributions can be summarized as below:
+Our contributions can be summarized as follows:
   • We establish the first theoretical analysis on why training $P(h|b)$ is intractable (combinatorial complexity).
@@
 O’Neill et al.
-(2025) attempt to model $P(h|b)$ directly; however, obtaining the reasoning trace towards $h$ from only $b$ via distillation is prohibitive (as discussed in § 6), and generating simulated reasoning traces directly from $b$ and the ground truth $h$ as input is very difficult to work (Wang et al., 2025).
-This work focuses exclusively on training LLMs for scientific discovery, distinct from related fields such as general scientific reasoning (Bai et al., 2025) or modeling of specific science modality (Jumper et al., 2021).
+(2025) attempt to model $P(h|b)$ directly; however, obtaining the reasoning trace towards $h$ from only $b$ via distillation is prohibitive (as discussed in § 7.1), and simulated reasoning traces generated directly from $b$ and the ground-truth $h$ are often difficult to make effective for training (Wang et al., 2025).
+This work focuses exclusively on LLMs for scientific discovery, distinct from related fields such as general scientific reasoning (Bai et al., 2025) or modeling of specific science modality (Jumper et al., 2021).
 3 Preliminary: The Decomposition Theory
@@
 This compositional view is well-supported by cognitive science literature (Koestler, 1964) and has proven empirically effective across scientific disciplines (Liu et al., 2025).
-For instance, the discovery of Backpropagation can be viewed as composing the background of “multilayer logistic regression” with the inspiration of the “chain rule” from calculus; similarly, a new synthesis method in chemistry might be derived by applying a specific “catalyst” (inspiration) to a known reaction mechanism.
+For instance, the discovery of Backpropagation can be viewed as composing the background of “multilayer logistic regression” with the inspiration of the “chain rule” from calculus; similarly, a new synthesis method in chemistry might be derived by applying a specific “catalyst” (inspiration) to a known reaction mechanism (background).
 Based on this formulation, we can approximate the intractable marginal likelihood $P(h\mid b)$ by decomposing the generation process into $k$ sequential steps.
@@
 Data Collection and Split.
-We collected 108,717 open-access papers from the NCBI database (Schoch et al., 2020), spanning biology, chemistry, and cognitive science.
+We collected 108,717 open-access papers from the NCBI database (Schoch et al., 2020), spanning biology, chemistry, medicine, medical imaging, psychology, and cognitive science.
 The corpus covers the period from January 2020 to October 2025.
@@
 For the IR task, the model is trained to identify the target inspiration via a generative selection process.
-The input consists of the query context and a candidate pool of 15 papers (1 positive and 14 negatives), each represented by its title and abstract.
-The negative set comprises easy negatives (randomly sampled papers published no later than the source paper) and hard negatives drawn from three specific categories: papers with keyword overlap with the source background, papers with keyword overlap with the ground-truth inspiration, and papers semantically similar to the ground-truth inspiration.
+The input consists of the query context and a pool of 15 papers (1 positive and 14 negatives), each represented by its title and abstract.
+The negative set comprises easy negatives (randomly sampled papers published no later than the source paper) and hard negatives drawn from three categories: papers with keyword overlap with the source background, papers with keyword overlap with the ground-truth inspiration, and papers semantically similar to the ground-truth inspiration.
 We acknowledge that these negatives might occasionally indicate valid alternative inspirations.
-However, the ground-truth citation represents the author’s explicit acknowledgment of causal influence, providing the most reliable supervision signal for prioritizing optimal inspirations over merely related ones.
+However, the ground-truth citation represents the author’s explicit acknowledgment of causal influence, providing the most reliable supervision for prioritizing optimal inspirations over merely related ones.
 Furthermore, this single-selection objective minimizes the penalty for potentially valid unlabelled candidates compared to independently labeling each candidate.
@@
 The model is conditioned on the research background $b$, the retrieved ground-truth inspiration $i_{j}$, and $h_{j-1}$, represented by the concatenation of $\{\Delta h_{0},\ldots,\Delta h_{j-1}\}$ (if $j\geq 1$).
-To enable effective rejection sampling, we developed a rubric-based evaluator to compare the quality of sampled hypotheses against the ground-truth $\Delta h_{j}$, ensuring that only high-quality samples are retained for training.
-The evaluation rubric is provided in § C.
+To enable effective rejection sampling, we developed a rubric-based evaluator to compare the quality of sampled hypotheses against the ground-truth $\Delta h_{j}^{*}$, ensuring that only high-quality samples are retained for training.
+The evaluation rubric is provided in § E.
 5.3 Method $\mathrm{II}$: Bounded Composition
@@
 We then employ the teacher model to generate hypotheses conditioned on these proxy inspirations rather than the exact ground truth.
-The generated outputs are rigorously evaluated against the quality control rubric described in § C.
+The generated outputs are rigorously evaluated against the quality control rubric described in § E.
 Crucially, for instances where proxies from multiple tiers yield passing hypotheses, we prioritize the sample from the most difficult tier (i.e., lowest similarity) for the final training set.
@@
 Despite the reduction to $O(N/M)$, a linear scan over the global knowledge base $\mathcal{I}$ remains inefficient.
-To enable logarithmic-time access (at least in the best-case scenario), we propose replacing the flat scan with a Hierarchical Best-First Search.
-This design draws inspiration from MOOSE-Chem2 (Yang et al., 2025a), which formulates fine-grained hypothesis discovery as a hierarchical search over varying levels of detail abstraction.
-Here, we adapt this concept by constructing the hierarchy directly over the inspiration space $\mathcal{I}$.
-By organizing $\mathcal{I}$ into a semantic search tree, we can rapidly navigate to the most promising subspaces, pruning irrelevant branches early in the process.
+To enable logarithmic-time search (at least in the best-case scenario), we replace the flat scan with a Hierarchical Best-First Search.
+This design is inspired by MOOSE-Chem2 (Yang et al., 2025a), which formally introduces the fine-grained hypothesis discovery task, formulates it as an optimization problem, and demonstrates that this optimization process can be made more tractable by organizing it hierarchically over multiple abstraction levels.
+While MOOSE-Chem2 develops this principle in the hypothesis space, here we adapt it to inspiration retrieval by constructing the hierarchy directly over the inspiration space $\mathcal{I}$.
 Tree Construction (Offline).
@@
 Best-First Search (Online).
-During inference, we traverse the tree using a Best-First Search (BFS) strategy.
+During inference, we traverse the tree using a Best-First Search strategy.
 We maintain a priority queue of candidate nodes, ranked by a path confidence score.
@@
 Model Accuracy
+Frontier Models
+GPT-4o-mini 41.99%
+Claude-Sonnet-4.6 45.02%
+DeepSeek-R1 45.11%
+GPT-4o 48.37%
+Gemini-3-Flash 51.44%
+GPT-5.4 51.50%
+Gemini-3-Pro 54.89%
+Baselines
 Random Selection 6.70%
@@
 MS-IR-7B 54.37%
+MS-7B 54.34%
 Table 2: Hypothesis composition performance given the ground-truth inspiration $i^{*}$.
@@
 Model Total Mot Mec Met Length
-R1-Distilled-Qwen-7B 4.34 1.96 1.40 0.97 231.02
-MS-HC-7B 5.08 2.21 1.58 1.29 204.12
-   w/ $1\times$ bounded 5.16 2.24 1.60 1.31 203.84
-   w/ $2\times$ bounded 5.15 2.25 1.58 1.32 205.17
-Inspiration Retrieval (IR).
-We fine-tuned R1-Distilled-Qwen-7B on a dataset of 152k retrieval examples, resulting in MS-IR-7B (where MS stands for MOOSE-Star).
-As shown in Table 1, this specialized training yields a substantial performance gain, boosting accuracy from 28.42% to 54.37%.
-Hypothesis Composition (HC).
-We fine-tuned R1-Distilled-Qwen-7B on 97k samples generated via teacher-based rejection sampling, resulting in MS-HC-7B (MS stands for MOOSE-Star).
-Performance is evaluated using the M3 score (detailed in § C), an LLM-based rubric summing scores for Motivation, Mechanism, and Methodology.
-Each dimension ranges from [0, 4] (representing 0%–100% information recall), yielding a total score of 12.
-Table 2 presents the results on ground-truth inputs.
-The fine-tuned MS-HC-7B significantly outperforms the baseline.
-Furthermore, results indicate that incorporating “bounded” training data (hypotheses generated from noisy inspirations) further enhances performance, suggesting that exposure to semantic variations during training improves reasoning robustness even when evaluating on perfect inspirations.
+Frontier Models
+GPT-4o-mini 5.18 2.26 1.64 1.28 175.29
+GPT-4o 5.99 2.43 1.95 1.60 264.93
+DeepSeek-R1 6.42 2.54 2.10 1.77 264.46
+Gemini-3-Flash 6.42 2.48 2.16 1.79 351.79
+Gemini-3-Pro 6.70 2.59 2.26 1.85 300.86
+Claude-Sonnet-4.6 7.28 2.78 2.41 2.08 510.26
+GPT-5.4 7.82 2.99 2.61 2.22 504.60
+Baselines
+R1-Distilled-Qwen-7B 4.05 1.96 1.30 0.80 231.02
+MS-HC-7B 4.68 2.13 1.46 1.09 204.12
+   w/ $1\times$ bounded 4.74 2.16 1.48 1.10 203.84
+   w/ $2\times$ bounded 4.73 2.15 1.48 1.09 205.17
+MS-7B 5.02 2.22 1.59 1.20 208.98
 Table 3: Hypothesis composition performance under varying levels of inspiration noise.
@@
 Model Total Mot Mec Met Total Mot Mec Met Total Mot Mec Met
-R1-Distilled-Qwen-7B 2.58 1.43 0.62 0.54 2.16 1.24 0.46 0.45 1.82 1.13 0.33 0.36
-MS-HC-7B 2.82 1.53 0.62 0.67 2.37 1.36 0.47 0.55 2.20 1.30 0.40 0.49
-   w/ $1\times$ bounded 2.96 1.55 0.70 0.71 2.52 1.41 0.53 0.58 2.19 1.25 0.44 0.49
-   w/ $2\times$ bounded 3.03 1.62 0.69 0.72 2.49 1.37 0.53 0.59 2.25 1.29 0.44 0.52
-6.2 Method $\mathrm{II}$: Bounded Composition
-Table 3 evaluates performance under noisy conditions, using proxy inspirations stratified by semantic similarity to $i^{*}$ (Easy, Medium, Hard).
-The results demonstrate that incorporating bounded training data significantly boosts M3 scores across all tiers compared to the base model.
-6.3 Method $\mathrm{III}$: Hierarchical Search
+R1-Distilled-Qwen-7B 2.72 1.50 0.70 0.51 2.27 1.35 0.55 0.38 2.00 1.24 0.46 0.31
+MS-HC-7B 3.10 1.63 0.81 0.66 2.59 1.46 0.61 0.53 2.39 1.38 0.54 0.48
+   w/ $1\times$ bounded 3.09 1.61 0.81 0.67 2.72 1.48 0.66 0.57 2.42 1.39 0.55 0.47
+   w/ $2\times$ bounded 3.18 1.67 0.83 0.68 2.74 1.51 0.66 0.57 2.56 1.45 0.58 0.53
+MS-7B 3.37 1.72 0.90 0.75 2.86 1.55 0.69 0.62 2.78 1.52 0.66 0.60
+Inspiration Retrieval (IR).
+We fine-tuned R1-Distilled-Qwen-7B on a dataset of 152k retrieval examples, resulting in MS-IR-7B (where MS stands for MOOSE-Star).
+As shown in Table 1, this specialized training yields a substantial performance gain, boosting accuracy from 28.42% to 54.37%.
+We verify in § C that this improvement reflects genuine retrieval ability rather than exploitation of superficial patterns.
+Hypothesis Composition (HC).
+We fine-tuned R1-Distilled-Qwen-7B on 97k samples generated via teacher-based rejection sampling, resulting in MS-HC-7B (MS stands for MOOSE-Star).
+Performance is evaluated using the M3 score (detailed in § E), an LLM-based rubric summing scores for Motivation, Mechanism, and Methodology.
+Each dimension ranges from [0, 4] (representing 0%–100% information recall), yielding a total score of 12.
+We use GPT-4o (OpenAI, 2024) as the judge model.
+Table 2 presents the results on ground-truth inputs.
+The fine-tuned MS-HC-7B significantly outperforms the baseline.
+Furthermore, results indicate that incorporating “bounded” training data (hypotheses generated from noisy inspirations) further enhances performance, suggesting that exposure to semantic variations during training improves reasoning robustness even when evaluating on perfect inspirations.
+Overall.
+These results suggest an interesting asymmetry between the two subtasks.
+For inspiration retrieval, targeted post-training appears sufficient to elevate a 7B model to near-frontier performance.
+By contrast, although hypothesis composition improves substantially over the base model, it remains clearly below frontier LLMs.
+This pattern suggests that retrieval may rely more on activating and organizing latent scientific associations already present in the base model, whereas high-quality composition likely depends more strongly on the base model’s underlying scientific reasoning and generation capacity.
+Motivated by the complementary nature of the two tasks, we jointly train on both IR and HC data (w/ $1\times$ bounded), yielding a unified model MS-7B.
+As shown in Tables 1 and 2, MS-7B retains IR accuracy (54.34% vs.
+54.37%) while improving HC performance (5.02 vs.
+4.68), suggesting that multi-task training provides a beneficial inductive bias for hypothesis composition without degrading retrieval.
 Table 4: Efficiency and accuracy of hierarchical search variants compared to a tournament baseline.
@@
    w/ motivation (detailed) 63.05 742.50
+6.2 Method $\mathrm{II}$: Bounded Composition
+Table 3 evaluates performance under noisy conditions, using proxy inspirations stratified by semantic similarity to $i^{*}$ (Easy, Medium, Hard).
+The results demonstrate that incorporating bounded training data significantly boosts M3 scores across all tiers compared to the base model.
+MS-7B further improves across all tiers, with the largest gain on Hard proxies (+0.22 total over w/ $2\times$ bounded), indicating that multi-task training strengthens robustness to retrieval noise.
+Similarly, GPT-4o is used as the judge model.
+6.3 Method $\mathrm{III}$: Hierarchical Search
 We evaluate search efficiency on a corpus of 3,035 ground-truth inspirations derived from the 1,658 test papers (October 2025).
@@
 The detailed variant achieves the best performance, confirming that more detailed and precise planning can reduce more computational effort required to locate $i^{*}$.
-7 Scaling Analysis
-7.1 Scaling Brute-force Sampling $P(h|b)$
 Table 5: Comparison of sampling pass rates between Brute-force (End-to-End) and Hypothesis Composition (Decomposed).
-While Brute-force performance collapses to 0.00% as complexity ($k$) increases, the decomposed HC module maintains a robust 47.33% pass rate, unlocking the data engine for RFT.
+Both generation and rubrics-based quality scoring are performed by R1-Distilled-Qwen-32B.
+While Brute-force collapses to 0.00% as complexity ($k$) increases, the decomposed HC module maintains a stable $\sim$47–52% per-step pass rate regardless of $k$, unlocking the data engine for RFT.
 Method Steps ($k$) Total Samples Passed Samples Pass Rate
-Brute-force $P(h|b)$ 1 31,314 653 2.09%
+Brute-force $P(h\mid b)$ 1 31,314 653 2.09%
 2 67,000 90 0.13%
@@
 Total 110,314 743 0.67%
-HC $P(\Delta h|i,b)$ 1 (per step) 195,301 92,427 47.33%
+[…]
+[…]
+[…]
+Total 195,301 96,879 49.60%
+7 Scaling Analysis
+7.1 Scaling Brute-force Sampling $P(h|b)$
 Direct distillation from the end-to-end distribution $P(h|b)$ is intractable due to the extreme scarcity of valid reasoning traces.
-Table 5 illustrates this bottleneck: finding a qualified hypothesis is rare even for simple cases ($k=1$), and becomes statistically impossible for multi-inspiration trajectories ($k\geq 2$) due to combinatorial explosion.
-In contrast, the decomposed HC step demonstrates a high success rate (47.33%).
+Table 5 illustrates this bottleneck: using R1-Distilled-Qwen-32B as both the generator and the rubric-based scorer, finding a qualified hypothesis is rare even for simple cases ($k=1$), and becomes statistically impossible for multi-inspiration trajectories ($k\geq 2$) due to combinatorial explosion.
+In contrast, the decomposed HC step demonstrates a high success rate (49.60%).
 This performance gain is driven by two structural advantages.
@@
 Evaluation Standard and The “Training Deadlock”.
-To understand why this performance gap determines training feasibility, we must examine the criteria used for harvesting RFT training data (detailed in § C).
-Both strategies use the same rubric, evaluating a generated hypothesis component against the ground truth across Motivation, Mechanism, and Methodology.
-To serve as a positive sample for RFT, we impose a filtering threshold: the generation must achieve a Total Score $\geq 8$ (out of 12).
-The structural difference lies in how this threshold applies to the two paradigms:
-  • Brute-force: Since the baseline generates the hypothesis $h$ holistically, a valid sample for a paper with $k\geq 2$ must successfully meet the filtering threshold for every constituent innovation $\Delta h_{j}^{*}$ ($j=1\dots k$) simultaneously.
-This strict conjunctive constraint drives the success rate to near-zero, creating a “training deadlock”: the model cannot generate enough valid samples to initiate RFT, rendering the training process impossible.
-  • HC: In contrast, the HC module is trained to synthesize a single increment $\Delta h_{j}$.
-Crucially, leveraging the decomposed data construction, we can condition the model on the ground-truth inspiration $i_{j}^{*}$.
-Under this setup, the model only needs to meet the threshold for the corresponding single $\Delta h_{j}^{*}$.
-Consequently, this efficient sampling capability breaks the training deadlock, making it feasible to train HC using RFT.
+Both strategies are scored by the same rubric (§ E), which evaluates Motivation, Mechanism, and Methodology against the ground truth.
+A generation must achieve a Total Score $\geq 8$ (out of 12) to qualify as a positive sample for RFT.
+  • Brute-force: Since the baseline generates the hypothesis $h$ holistically, a valid sample for a paper with $k$ steps must meet the threshold for every constituent innovation $\Delta h_{j}^{*}$ ($j=1\dots k$) simultaneously.
+As $k$ increases, the likelihood of satisfying all $k$ criteria jointly vanishes, creating a “training deadlock” where the model cannot generate enough valid samples to initiate RFT.
+  • HC: In contrast, the HC module synthesizes a single increment $\Delta h_{j}$ conditioned on the corresponding inspiration $i_{j}^{*}$.
+The reduced scope and provided guidance together yield a stable $\sim$50% pass rate, breaking the training deadlock and making RFT feasible.
 7.2 Scaling Training of MOOSE-Star
-We analyze the scaling behavior of the IR and HC models with training data ranging from $10^{0}$ to $10^{5}$ samples (Figure 2(a) and 2(b)).
+We analyze the scaling behavior of the dedicated IR and HC models (i.e., MS-IR-7B and MS-HC-7B, each trained on its respective subtask data) with training data ranging from $10^{0}$ to $10^{5}$ samples (Figure 2(a) and 2(b)).
 The IR model demonstrates a standard log-linear improvement in accuracy.
@@
 7.3 Test-time Scaling of MOOSE-Star
-Figure 3(a) compares the test-time scaling efficiency of MOOSE-Star against the brute-force baseline.
+Figure 3(a) compares the test-time scaling efficiency of the MOOSE-Star system against the brute-force baseline.
 The evaluation is conducted on a subset of TOMATO-Star comprising the first 109 test cases.
@@
 MOOSE-Star achieves steady progress, continuously solving more complex cases as the search budget increases, eventually reaching 100% coverage of the test set at $\sim$6,000 inference calls.
-Conversely, the brute-force baseline saturates at $\sim$41.3%.
+Conversely, the brute-force baseline saturates at $\sim$38.5%.
 Figure 3(b) provides a breakdown of this saturation, revealing the steep “complexity cliff” where unstructured sampling fails.
@@
 However, this volume advantage is rapidly negated by the combinatorial explosion.
-As the reasoning chain extends, the baseline’s performance collapses—plummeting to $\sim$36% for $k=2$ and further evaporating to a negligible $\sim$8% for $k=3$.
+As the reasoning chain extends, the baseline’s performance collapses—plummeting to $\sim$37% for $k=2$ and further evaporating to a negligible $\sim$8% for $k=3$.
 This dramatic decay empirically confirms that without the structural guidance of decomposition, even nearly ten thousand unguided attempts are statistically insufficient to break the combinatorial complexity barrier.
+We provide an additional analysis on temporally held-out October 2025 papers in § D, including both quantitative joint retrieval–composition evaluation and a case study.
 8 Conclusion
@@
 Figure 4: MOOSE-Star Concept Figure
+The name MOOSE-Star reflects the intuition behind our framework: discovering the right inspiration from a massive scientific corpus is like finding a specific star in a vast universe.
+The challenge is not only the size of the search space, but also how to navigate it efficiently.
 Appendix B Theoretical Derivation: Rigorous Decomposition of $P(h\mid b)$ via Hierarchical MDP
@@
 where $h_{0}^{(\pi)}=\emptyset$, $h_{k}^{(\pi)}=h$, and $h_{j}^{(\pi)}$ denotes the intermediate hypothesis state at step $j$ under the specific permutation $\pi$.
-Appendix C Rubrics for Comparing Generated Hypothesis with Groundtruth Hypothesis
+Appendix C Ablation of Inspiration Retrieval Experiment: Genuine Retrieval vs. Distribution Shortcuts
+A potential concern with the inspiration retrieval (IR) evaluation (Table 1) is that the model may exploit distributional patterns in the candidate pool rather than performing genuine content-based retrieval.
+In the standard setting, each sample contains 1 ground-truth inspiration and 14 negatives, where $\sim$35% of negatives are hard (keyword-overlap or embedding-similar to the ground truth) and $\sim$65% are randomly sampled.
+Since training and test sets use the same negative construction pipeline, the model could potentially learn a shortcut: detecting the cluster of semantically related papers and selecting from it, without truly understanding the research context.
+To investigate this, we construct three ablation test sets by modifying only the negative candidates while keeping the ground truth and query context identical:
+  • All Random: All 14 negatives are randomly sampled (no hard negatives).
+Without a cluster of related papers in the pool, a model that relied on cluster detection loses its signal, while a model doing genuine content matching benefits from the easier distractors.
+  • All Hard: All 14 negatives are hard negatives (keyword/embedding similar to the ground truth). This is the most challenging setting, removing easy negatives that can be trivially eliminated.
+  • Decoy Cluster: In the original setting, the hard negatives are semantically related to the ground-truth inspiration, forming a cluster of topically similar papers around the correct answer.
+To test whether models simply detect this cluster, we replace these hard negatives with those harvested for a different test sample’s ground-truth inspiration (the “decoy”).
+This produces a cluster of papers that are semantically coherent with each other—but irrelevant to the current query’s research context and correct answer.
+The remaining slots are filled with random papers, preserving the same hard/random ratio ($\sim$35/65%) as the original.
+A model that selects by detecting “which group of papers looks related to each other” will be drawn to the decoy cluster; a model that matches candidates against the query context will ignore it.
+Table 6 shows the results. The “Original” column is the standard setting from Table 1. Three consistent patterns emerge:
+(1) All Random $\gg$ Original: All models show substantially higher accuracy (e.g., MS-IR-7B: 54.37% $\to$ 78.58%), confirming that hard negatives are the primary source of task difficulty.
+(2) All Hard $\approx$ Original: Accuracy remains similar (e.g., MS-IR-7B: 54.37% vs.
+53.42%), indicating that the original $\sim$35% hard negative ratio already captures most of the difficulty; adding more hard negatives has diminishing effect.
+(3) Decoy Cluster $\approx$ All Random: Accuracies under Decoy Cluster closely match All Random across all models, demonstrating that retrieval decisions are based on content matching between the query context and candidate papers, not on detecting clusters of semantically related candidates.
+These results confirm that MS-IR-7B’s strong performance on the original benchmark reflects genuine inspiration retrieval ability acquired through SFT training, not exploitation of distributional artifacts in the negative construction.
+MS-7B exhibits the same pattern, confirming that multi-task training preserves this genuine retrieval ability.
+Table 6: IR ablation study: accuracy (%) under modified negative construction.
+Model Original All Random All Hard Decoy Cluster
+Frontier Models
+GPT-4o-mini 41.99 77.43 41.00 77.43
+Claude-Sonnet-4.6 45.02 91.44 58.69 92.38
+DeepSeek-R1 45.11 75.29 43.84 75.64
+GPT-4o 48.37 82.18 49.43 82.02
+Gemini-3-Flash 51.44 78.86 48.05 80.84
+GPT-5.4 51.50 78.93 51.34 80.97
+Gemini-3-Pro 54.89 84.71 54.66 85.60
+Baselines
+Random Selection 6.70 6.70 6.70 6.70
+R1-Distilled-Qwen-7B 28.42 45.27 25.30 42.71
+MS-IR-7B 54.37 78.58 53.42 77.65
+MS-7B 54.34 79.25 52.52 78.38
+Appendix D Temporal Generalization on Temporally Held-Out Papers
+To complement the test-time scaling analysis in § 7.3, we further examine whether MOOSE-Star can operate on genuinely temporally held-out scientific papers.
+Concretely, all evaluations in this section are conducted on papers published in October 2025, which are excluded from both the base model’s pre-training period and our post-training data.
+This setting allows us to assess whether the framework can retrieve key inspirations and use them to construct high-quality hypotheses for future papers whose scientific content is unseen during both pre-training and post-training.
+The goal of this section is not to claim prospective real-world validation, which requires waiting for future publications, but rather to provide controlled temporal out-of-distribution (OOD) evidence under a strict chronological split.
+We first present quantitative results on held-out future papers, measuring retrieval success and downstream hypothesis composition quality.
+We then analyze whether the main bottleneck lies in retrieval or composition.
+Finally, we provide a full $k>1$ case study showing how MOOSE-Star retrieves two ground-truth inspirations from the inspiration corpus and composes a two-step hypothesis trajectory that closely aligns with the actual paper.
+Table 7: ”Retrieval success” denotes the fraction of all 200 test cases whose ground-truth inspiration is retrieved within Rank $\leq K$.
+Other entries report the fraction of all 200 test cases for which this retrieval condition holds and the HC module, when conditioned on the ground-truth inspiration, reaches the corresponding score threshold.
+Rank HC Score MS-HC-7B MS-7B GPT-4o GPT-5.4 DeepSeek-R1 Claude-Sonnet-4.6 Gemini-3-Pro
+$\leq 25$ $\geq 4$ 14.0% 17.0% 19.0% 19.0% 18.5% 18.5% 18.5%
+$\geq 5$ 10.0% 14.0% 18.0% 18.5% 17.0% 18.5% 18.0%
+$\geq 6$ 8.0% 10.0% 14.5% 16.5% 14.0% 16.0% 16.0%
+$\geq 7$ 2.0% 4.5% 10.5% 16.0% 13.0% 16.0% 14.5%
+$\geq 8$ 0.0% 0.5% 5.5% 12.5% 7.5% 11.0% 8.5%
+$\geq 9$ 0.0% 0.0% 2.0% 8.0% 3.0% 4.0% 4.0%
+$\geq 10$ 0.0% 0.0% 1.0% 7.0% 1.5% 2.5% 1.5%
+$\geq 11$ 0.0% 0.0% 0.0% 3.5% 0.0% 0.0% 0.5%
+$\geq 12$ 0.0% 0.0% 0.0% 1.5% 0.0% 0.0% 0.0%
+Retrieval success 20.0%
+$\leq 50$ $\geq 4$ 20.0% 22.5% 26.0% 25.5% 25.0% 25.0% 25.0%
+$\geq 5$ 15.0% 17.5% 23.5% 25.0% 23.5% 25.0% 24.0%
+$\geq 6$ 12.5% 13.0% 19.0% 22.5% 20.5% 22.0% 21.5%
+$\geq 7$ 4.0% 5.0% 13.5% 20.5% 18.0% 21.0% 19.5%
+$\geq 8$ 0.5% 0.5% 6.5% 16.0% 10.5% 13.0% 11.5%
+$\geq 9$ 0.0% 0.0% 2.5% 10.5% 5.0% 4.5% 5.5%
+$\geq 10$ 0.0% 0.0% 1.5% 8.5% 2.0% 3.0% 1.5%
+$\geq 11$ 0.0% 0.0% 0.5% 4.0% 0.0% 0.5% 0.5%
+$\geq 12$ 0.0% 0.0% 0.0% 2.0% 0.0% 0.5% 0.0%
+Retrieval success 28.0%
+$\leq 100$ $\geq 4$ 20.5% 23.5% 27.0% 26.5% 26.0% 26.0% 26.0%
+$\geq 5$ 15.5% 17.5% 24.0% 26.0% 24.5% 26.0% 25.0%
+$\geq 6$ 13.0% 13.0% 19.5% 23.0% 21.5% 23.0% 22.5%
+$\geq 7$ 4.0% 5.0% 14.0% 21.0% 19.0% 22.0% 20.5%
+$\geq 8$ 0.5% 0.5% 6.5% 16.5% 11.0% 13.0% 11.5%
+$\geq 9$ 0.0% 0.0% 2.5% 10.5% 5.5% 4.5% 5.5%
+$\geq 10$ 0.0% 0.0% 1.5% 8.5% 2.0% 3.0% 1.5%
+$\geq 11$ 0.0% 0.0% 0.5% 4.0% 0.0% 0.5% 0.5%
+$\geq 12$ 0.0% 0.0% 0.0% 2.0% 0.0% 0.5% 0.0%
+Retrieval success 29.0%
+D.1 Quantitative Evidence on Temporally Out-of-Distribution Data
+Our test set consists entirely of papers published in October 2025.
+Both MS-IR-7B and MS-HC-7B are derived from R1-Distilled-Qwen-7B and fine-tuned exclusively on pre-October 2025 data.
+We therefore regard this evaluation as a strict temporal OOD setting.
+We define a temporal reconstruction success using two criteria: (1) the hierarchical search ranks the ground-truth inspiration within the top-$K$ candidates, and (2) the HC module, when conditioned on the ground-truth inspiration, produces a hypothesis that reaches at least a target M3 score, as judged by GPT-4o.
+In this evaluation, retrieval is performed over an inspiration corpus containing 3,035 candidates.
+We evaluate this criterion on 200 test samples using MS-IR-7B for retrieval and several hypothesis composition backends.
+In addition to the single-task MS-HC-7B, we include MS-7B, a multitask variant jointly trained on both IR and HC (w/ $1\times$ bounded) data, which consistently outperforms the single-task model (e.g., $17.5\%$ vs.
+$15.0\%$ at Rank $\leq 50$, M3 $\geq 5$).
+Table 7 reports the resulting success rates under three retrieval thresholds: Rank $\leq 25$, Rank $\leq 50$, and Rank $\leq 100$.
+These thresholds correspond to retrieving the correct inspiration within approximately the top $0.8\%$, $1.6\%$, and $3.3\%$ of the inspiration corpus, respectively.
+Several observations emerge from these results.
+First, retrieval appears to be the main bottleneck.
+Among the 200 temporally held-out test samples, the ground-truth inspiration is ranked within the top 25 for $20.0\%$ of cases, within the top 50 for $28.0\%$, and within the top 100 for only slightly more cases ($29.0\%$).
+Second, hypothesis composition quality still matters substantially once the correct inspiration has been identified.
+Under Rank $\leq 50$, MS-HC-7B achieves $4.0\%$ success at M3 $\geq 7$, while stronger frontier models used as HC backends achieve substantially higher rates ($13.5\%$$21.0\%$).
+This gap indicates that stronger composition models can make more effective use of the scientific signal once the relevant inspiration is available.
+D.2 Case Study: A Two-Inspiration Discovery Reconstructed from a Pre-2025 Inspiration Corpus
+We next present a case study on a temporally unseen paper requiring $k=2$ inspirations.
+This example is informative because it exposes the full sequential structure of MOOSE-Star: the framework must retrieve one ground-truth inspiration, compose a meaningful intermediate hypothesis, and then retrieve and integrate a second ground-truth inspiration that extends the scientific trajectory.
+Unlike a $k=1$ example, this setting more directly illustrates how multiple inspirations can be combined into a coherent multi-step discovery process.
+Target paper.
+TREM2 Impedes Recovery After Spinal Cord Injury by Regulating Microglial Lysosomal Membrane Permeabilization-Mediated Autophagy (Cell Proliferation, October 2025; DOI: 10.1111/cpr.70047).
+Research question.
+How to enhance microglial survival and functional recovery after spinal cord injury by regulating lysosomal integrity and autophagy flux?
+Background survey.
+Spinal cord injury (SCI) involves irreversible primary mechanical damage followed by secondary injury phases.
+Microglia, the primary immune responders in the CNS, proliferate post-injury and their depletion exacerbates damage.
+Protecting microglial numbers is a known therapeutic strategy, but mechanisms regulating their survival/function remain unclear.
+TREM2 regulates microglial activation states (e.g., DAM phenotype) in neurodegenerative diseases but has not been studied in SCI.
+Single-cell data suggest a link between TREM2 and lysosomal pathways in microglia.
+Lysosomal dysfunction (reduced activity, impaired autophagy) contributes to neuronal death in SCI, but its role in microglial survival is unexplored.
+Autophagy maintains cellular homeostasis but requires functional lysosomes.
+Retrieved inspirations.
+The two retrieved ground-truth inspirations by MS-IR-7B are shown in Table 8.
+Table 8: Retrieved ground-truth inspirations for the temporally held-out TREM2 case. For this $k=2$ example, MOOSE-Star ranks the two ground-truth inspirations at 12/3035 and 1/3035, respectively.
+Step Ground-Truth Inspiration DOI Retrieval Rank
+0 Underlying Mechanism of Lysosomal Membrane Permeabilization in CNS Injury: A Literature Review (Molecular Neurobiology, 2024) 10.1007/s12035-024-04290-6 12 / 3035
+1 TFEB links autophagy to lysosomal biogenesis (Settembre et al., Science, 2011) 10.1126/science.1204592 1 / 3035
+Step 0: Generated Hypothesis (GPT-5.4 as HC model; M3 = 9/12).
+Inspiration: Lysosomal membrane permeabilization (LMP) is the key upstream lysosomal lesion in SCI microglia, and stabilizing lysosomal membranes to prevent cathepsin leakage will preserve autophagy flux and microglial survival.
+- Motivation (WHY): Current SCI approaches describe lysosomal dysfunction and impaired autophagy broadly, but they do not specify the proximal lysosomal failure that converts injury stress into microglial death and loss of function.
+The inspiration paper identifies LMP as a central event in CNS injury that links oxidative/inflammatory stress to release of lysosomal hydrolases, programmed cell death, and autophagy disruption.
+Adapting this to SCI suggests that microglial loss may be driven not simply by reduced lysosome activity, but by physical destabilization of lysosomal membranes.
+Targeting LMP therefore fills the mechanistic gap between SCI-induced stress and both outcomes of interest: diminished microglial survival and defective autophagic clearance.
+- Mechanism (HOW IT WORKS): After SCI, microglia are exposed to ROS, lipid peroxidation, excitotoxic/calcium stress, and inflammatory mediators that can permeabilize lysosomal membranes.
+LMP causes partial or extensive escape of cathepsins and other lysosomal contents into the cytosol, which then triggers lysosome-dependent cell death and amplifies apoptosis/pyroptotic-inflammatory signaling.
+At the same time, membrane-compromised lysosomes lose degradative competence and acidity, so autophagosomes can no longer be efficiently processed to completion, producing stalled autophagy flux rather than protective recycling.
+Therefore, preserving lysosomal membrane integrity should simultaneously: (i) retain cathepsins within lysosomes, (ii) maintain lysosomal acidification and enzyme compartmentalization, (iii) allow autophagosome-lysosome degradation to proceed, and (iv) reduce microglial death and dysfunctional inflammatory activation.
+The net effect should be better preservation of viable, functional microglia after SCI.
+- Methodology (HOW IT’S INTEGRATED): In SCI microglia, explicitly model and measure LMP as the central lysosomal integrity variable rather than only total lysosome abundance or generic autophagy markers.
+Induce SCI or SCI-like stress in vivo/in vitro, then quantify LMP using lysosomal rupture/permeabilization readouts such as galectin puncta recruitment, cytosolic cathepsin B/D/L redistribution, lysosomal pH loss, acridine orange or LysoTracker leakage, and ultrastructural membrane damage.
+In parallel, measure autophagy flux with LC3-II turnover and p62/SQSTM1 clearance under lysosomal blockade controls, and relate these to microglial survival and phenotype.
+Test whether lysosome-preserving interventions—such as lysosomal membrane stabilizers, cathepsin release/activity suppression, or anti-oxidative strategies that specifically reduce membrane lipid damage—rescue microglia by lowering LMP.
+Mechanistic validation should require showing that interventions improve survival and functional recovery only when they reduce LMP, retain lysosomal cathepsins, and restore degradative autophagy flux.
+This establishes LMP prevention as the actionable lysosomal-integrity strategy for enhancing microglial survival after SCI.
+Step 0: Ground-Truth Hypothesis.
+Inspiration: Lysosomal membrane permeabilization (LMP) causing cell death
+- Motivation (WHY): The specific limitation addressed is the lack of understanding of how lysosomal dysfunction contributes to microglial death after spinal cord injury (SCI).
+Existing methods focus on protecting microglial numbers but fail to address the molecular mechanisms of microglial survival post-injury.
+LMP, a process where lysosomal membranes lose integrity, has been implicated in cell death in non-SCI contexts (e.g., cancer, neurodegeneration), but its role in microglial death after SCI remains unexplored.
+This gap is critical because lysosomal integrity is essential for maintaining autophagy, a key survival mechanism for stressed cells.
+- Mechanism (HOW IT WORKS): LMP occurs when lysosomal membranes lose their integrity, leading to the leakage of lysosomal enzymes (e.g., cathepsins) into the cytoplasm.
+This leakage triggers a cascade of events, including the activation of pro-apoptotic proteins (e.g., galectin-3, Gal3) and the loss of lysosomal function, ultimately resulting in cell death.
+In the context of SCI, this mechanism is particularly relevant because microglia are exposed to severe oxidative stress and inflammatory mediators, which can compromise lysosomal membranes.
+The causal chain is as follows: (1) SCI induces oxidative stress and inflammation.
+(2) Oxidative stress disrupts lysosomal membrane integrity.
+(3) LMP leads to cathepsin leakage and Gal3 activation.
+(4) Gal3 activation triggers pro-apoptotic signaling pathways.
+(5) Microglial death exacerbates secondary injury.
+[…]
+Step 1: Generated Hypothesis (GPT-5.4 as HC model; M3 = 12/12).
+Inspiration: TFEB-driven coupling of lysosomal biogenesis to autophagy flux restoration
+- Motivation (WHY): The previous hypothesis identifies lysosomal membrane damage as a cause of microglial death after spinal cord injury (SCI), but it does not explain how surviving microglia could rebuild degradative capacity once lysosomes are injured or insufficient.
+A key remaining limitation is that lysosomal integrity and autophagy flux are still treated mainly as damage readouts rather than as a coordinated repair program.
+The inspiration paper suggests that TFEB is a master regulator that simultaneously induces lysosomal biogenesis and the full autophagy machinery.
+This is directly relevant because microglia in SCI must clear damaged organelles, myelin debris, and inflammatory cargo under intense stress; without a transcriptional program that replenishes lysosomes and reopens autophagic flux, microglia would remain vulnerable to persistent lysosomal failure and cell death.
+Thus, the specific contribution of this inspiration is the hypothesis that impaired TFEB activation is an upstream reason why SCI microglia fail to restore lysosomal competence, and that enhancing TFEB activity can rescue microglial survival and function.
+- Mechanism (HOW IT WORKS): After SCI, microglia experience oxidative, inflammatory, and metabolic stress that increases degradative demand while damaging lysosomes.
+In this setting, TFEB functions as a transcriptional switch that can restore the lysosomal-autophagy system if it translocates to the nucleus.
+Once activated, TFEB is expected to induce a coordinated gene program covering: (1) lysosomal biogenesis and acidification (e.g., LAMP1, lysosomal hydrolases, v-ATPase-related genes), (2) autophagosome formation machinery, (3) autophagosome-lysosome fusion components, and (4) substrate degradation capacity.
+In SCI microglia, defective TFEB nuclear localization or persistent inhibitory phosphorylation would prevent this adaptive program, producing too few functional lysosomes, incomplete autophagic degradation, p62/LC3 cargo accumulation, and reduced capacity to clear damaged material.
+This would amplify lysosomal stress and sensitize cells to the lysosomal injury pathway defined in the previous hypothesis.
+In contrast, TFEB activation should replenish the lysosome pool, improve lysosomal enzyme content and acidification, increase fusion of autophagosomes with lysosomes, and re-establish autophagy flux.
+Functionally, this should enhance microglial survival, preserve phagocytic clearance, and support a reparative response after SCI.
+Because TREM2 is already linked by background data to lysosomal pathways, an additional testable extension is that TREM2 may promote microglial resilience partly by permitting TFEB activation/nuclear translocation; therefore, TFEB dysfunction may be more severe in Trem2-deficient microglia.
+- Methodology (HOW IT’S INTEGRATED): To integrate this concept into the existing SCI-microglia framework, TFEB is introduced as an upstream regulatory node controlling lysosomal recovery and autophagy completion: (1) Determine whether TFEB activation is impaired after SCI and in Trem2 deficiency—use wild-type and Trem2-/- SCI mice, measure TFEB subcellular localization by immunofluorescence and nuclear/cytoplasmic fractionation, quantify TFEB phosphorylation status, and perform qPCR/RNA-seq for TFEB target genes.
+(2) Assess whether TFEB status predicts restoration of lysosomal function—measure lysosome abundance using LAMP1/LAMP2 staining, lysosomal acidity using LysoTracker/LysoSensor, and compare between WT and Trem2-/- microglia.
+(3) Directly test whether TFEB restores autophagy flux—evaluate LC3-II and p62/SQSTM1 with and without lysosomal blockade (e.g., bafilomycin A1), use tandem fluorescent LC3 reporter (mCherry-GFP-LC3) to quantify autophagosome maturation.
+(4) Manipulate TFEB causally—activate TFEB genetically or pharmacologically, inhibit via siRNA/shRNA or CRISPR interference, test rescue of lysosomal biogenesis, autophagy flux, and cell survival, especially in Trem2-/- microglia.
+(5) Link TFEB-mediated restoration to microglial functional recovery—measure viability, phagocytosis, and inflammatory gene expression; in vivo, test whether microglia-targeted TFEB activation improves tissue sparing and behavioral recovery.
+(6) Integrate with the prior LMP framework—use lysosomal damage assays as downstream comparators to show that TFEB activation reduces susceptibility to lysosomal failure by increasing lysosome renewal and autophagy completion.
+Step 1: Ground-Truth Hypothesis.
+Inspiration: TFEB-mediated lysosomal biogenesis under cellular stress
+- Motivation (WHY): The specific limitation addressed is the lack of understanding of how to enhance autophagy in microglia after SCI.
+While autophagy is recognized as a critical survival mechanism, its regulation in the context of lysosomal dysfunction remains unclear.
+Existing methods do not address how to activate lysosomal biogenesis in stressed microglia, which is essential for maintaining autophagic flux and preventing cell death.
+- Mechanism (HOW IT WORKS): Transcription factor EB (TFEB) is a master regulator of lysosomal biogenesis and autophagy.
+Under cellular stress, TFEB translocates from the cytoplasm to the nucleus, where it activates the transcription of genes encoding lysosomal enzymes (e.g., CTSD) and autophagy-related proteins (e.g., LC3).
+This mechanism ensures that cells can degrade damaged organelles and recycle nutrients, promoting survival under stress conditions.
+In the context of SCI, this mechanism is particularly relevant because: (1) Microglia are exposed to oxidative stress and inflammation post-injury.
+(2) Oxidative stress disrupts lysosomal integrity, impairing autophagy.
+(3) TFEB activation can counteract this by promoting lysosomal biogenesis and restoring autophagy.
+- Methodology (HOW IT’S INTEGRATED): To investigate TFEB-mediated lysosomal biogenesis in microglia after SCI: (1) Experimental design: Wild-type (Wt) and Trem2-/- mice are subjected to SCI.
+Microglia are isolated from the spinal cord tissue 7 days post-injury.
+(2) TFEB activation: In vitro experiments are performed using BV2 cells treated with L-Leucyl-L-Leucine methyl ester (LLOMe) to induce LMP.
+Small interfering RNA (siRNA) targeting TREM2 is used to silence its expression.
+(3) Marker assessment: Immunofluorescence staining is performed to measure TFEB nuclear translocation.
+Western blotting is used to quantify the levels of TFEB, p-Syk, and autophagy markers (e.g., LC3-II, P62) in nuclear and cytosolic fractions.
+(4) Quantification: The mean fluorescence intensity of nuclear TFEB is measured using ImageJ software.
+The ratio of LC3-II to P62 is calculated to assess autophagic flux.
+Key Integration Between Inspirations: The two inspirations are integrated through the Syk signaling pathway.
+Specifically: (1) TREM2 knockout reduces Syk phosphorylation, which inhibits Syk-mediated LMP induction.
+(2) Reduced Syk phosphorylation promotes TFEB nuclear translocation, activating lysosomal biogenesis and autophagy.
+(3) This dual mechanism (LMP repair + autophagy enhancement) rescues microglial survival and promotes functional recovery after SCI.
+Discussion.
+This case illustrates the full sequential structure of MOOSE-Star on a temporally unseen paper.
+The model input (research question plus background survey) contains no mention of LMP, TFEB, or lysosomal biogenesis; these key scientific concepts enter the reasoning process only through the retrieved inspirations.
+Moreover, the two retrieved inspirations do not function as isolated fragments.
+Step 0 identifies LMP as the proximal mechanism of lysosomal failure, while Step 1 introduces TFEB as a regulatory program for restoring lysosomal function and autophagy flux.
+Together, the two steps closely reconstruct the two-part hypothesis structure of the actual paper.
+Appendix E Rubrics for Comparing Generated Hypothesis with Groundtruth Hypothesis
 Scoring Rubric (0–4 for each dimension)